# MIPHEI-vit: Multiplex Immunofluorescence Prediction from H&E Images using ViT Foundation Models

### v1

@@ 3.4 Data Preprocessing
We designed a preprocessing pipeline tailored to our H&E-to-mIF prediction task, addressing WSI registration, artifact removal, autofluorescence subtraction with normalization, and pseudo-label extraction, as illustrated in Figure 1.
@@ 3.4.1 H&E to mIF registration on consecutive cuts
Accurate registration between H&E and mIF images is crucial for training and evaluating our models. For Orion and HEMIT datasets, data was already registered Lin et al. (2023b); Bian et al. (2024). For the IMMUcan dataset, we used Valis Gatenbee et al. (2023) for registration of consecutive H&E and mIF slides.
@@ 3.4.2 Tile Selection
To reduce artifacts from restaining and acquisition in the ORION dataset, we filtered aligned H&E and mIF tiles using several quality control steps. We used thresholding on an empty mIF channel—without antibody staining but capturing shared noise—to identify artifacts affecting all mIF channels. Poor H&E quality tiles were identified by clustering H-optimus-0 embeddings and discarding clusters associated to obvious artifacts. We also manually annotated misaligned tiles and trained a CNN to detect them automatically from H&E and DAPI images. In total, about 40k tiles (10%) were excluded.
For the IMMUcan dataset, where H&E and mIF are from consecutive sections, we selected well-aligned tiles to ensure reliable correlation analysis based on three criteria: Pearson correlation of aligned nuclei density maps from 32×32 patches (threshold=0.25 per tile), tissue overlap ($IoU>0.5$), and tissue percentage ($>40\%$). We extracted $1024\times 1024$ pixel tiles at 20x (0.26 mm²), retaining 17k tiles covering 44.4 cm².
@@ 3.4.3 Autofluorescence Subtraction & Data Normalization
Autofluorescence (AF), captured as an independent channel $I_{AF}$, refers to light naturally emitted by the tissue across channels, introducing noise in other markers. We subtract the AF from each marker channel $I_{IF}^{c}$, using $\lambda^{c}$ and $b^{c}$ which were manually adjusted using a Napari Sofroniew et al. (2025) tool we developed:
$I_{cor}^{c}=\max(0,I_{IF}^{c}-\lambda^{c}\cdot I_{AF}+b^{c})$
Next, each channel is normalized using the 99.9th percentile $q_{0.999}^{c}$, computed per marker from the distribution of foreground pixel intensities across the training set, and then log-transformed:
$I_{norm}^{c}=255\cdot\log\left(\frac{\min(I_{cor}^{c},q_{0.999}^{c})}{q_{0.999}^{c}}+1\right)$
This logarithmic transformation compresses high-intensity values and reduces the impact of extreme outliers, while normalization ensures a consistent dynamic range across markers. The autofluorescence Napari tool, selected parameters, code, model, and data access instructions are available on our GitHub, within the Sanofi Public organization, under specific license conditions including a limitation to non-commercial use only.
@@ 3.4.4 Single-Cell Pseudo-Label Extraction
To establish a ground truth for evaluating our model’s ability to identify marker-positive cells from H&E images, we used a standard cell type calling approach on mIF images Lin et al. (2023a). We first segmented nuclei from the DAPI channel using Cellpose Stringer et al. (2021) fine-tuned on our data. As a proxy for cell regions, we dilated the nuclear regions by 2 $\mu m$. Single-cell analysis was performed by extracting mean fluorescence intensities, followed by unsupervised gating using a Gaussian Mixture Model (GMM) to distinguish positive from negative cells Zhang et al. (2022), with posterior probabilities estimated from the GMM. Although effective, this approach is sensitive to artifacts, approximate boundaries, and signal spillover. To improve robustness, we implemented a hierarchical gating strategy making sure that the biological marker hierarchy was preserved. For instance, given that CD3 positive cells are also CD45 positive, we kept CD3 positivity only for CD45 positive cells. By applying these rules, we obtained a high confidence annotation of our cells.
For the IMMUcan dataset, we used single-cell data provided by the consortium, which was generated using in-house clustering and nucleus segmentation. Since the same nuclei are not necessarily present in consecutive H&E and mIF sections, we applied Hoverfast Liakopoulos et al. (2024) for nucleus segmentation from H&E images, allowing us to perform the cell-level correlation analysis explained in section 5.2.
@@ 4 Methodology
This section outlines our overall approach for predicting mIF images from H&E images. We used the Orion dataset for training, and HEMIT and IMMUcan datasets for testing.
@@ 4.1 Model Architecture
Our model utilizes a U-Net generator to perform the image translation task from H&E to mIF signals (Figure 2). The architecture supports various encoder types, including CNNs like ConvNeXt v2 Woo et al. (2023), Swin Transformer Liu et al. (2021), and plain ViTs Dosovitskiy et al. (2020), enabling integration of recent foundation models in histology, such as CTransPath, Univ2, and H-optimus-0 Wang et al. (2022); Chen et al. (2024); Saillard et al. (2024). CNN-based encoders naturally align with the classical U-Net design, providing pyramidal features. For ViTs, which maintain fixed feature dimensions across layers, we explore two alternatives to generate multi-scale features: (1) convolutional transpose for upsampling token features, as used in UNETR Hatamizadeh et al. (2022) and CellViT Hörst et al. (2024), and (2) a hybrid approach inspired by ViTMatte Yao et al. (2024), where a convolutional stream extracts pyramidal features while ViT features serve as the bottleneck. Unlike original ViTMatte, which concatenates a trimap to the input image, we use only the H&E RGB image and omit convolutional necks and window attention. For simplicity, we still refer to this variant as VitMatte.
Building on these encoded features, the decoder reconstructs outputs using nearest interpolation for upsampling, combined with skip connections, dual 3×3 convolutional layers, batch normalization, and ReLU activation. Multiple output heads predict various mIF signals from the same decoder outputs, with a final Tanh activation.
@@ 4.2 Loss Function
We train MIPHEI using a weighted Mean Square Error (MSE) loss to ensure accurate translation from H&E to mIF signals while accounting for varying signal intensities and prevalence across markers. Each marker’s loss is scaled by the inverse of its standard deviation to balance contributions Mai et al. (2021).
@@
$\mathcal{L}_{\text{wMSE}}=\frac{\lambda}{M}\sum_{j}\frac{1}{\sigma_{j}^{2}}L_{\text{MSE},j}$
@@ 4.3 Foundation Model Training Strategies
While CNN-based encoders within the U-Net architecture are fully fine-tuned in all experiments, we explore two efficient training strategies for large ViT-based foundation model encoders: (1) decoder-only training with a frozen foundation encoder, leveraging robust pretrained features while reducing trainable parameters; (2) Low-Rank Adaptation (LoRA) Hu et al. (2022) with rank = 8 and $\alpha=1$, which adapts a pretrained ViT encoder by adding trainable low-rank matrices to the Query and Value projections of the attention layers.
@@ 5 Experiments
Figure 4: Prediction pipeline and prediction visualization: (a) Inference pipeline: mIF images are first generated using a trained U-Net model. Predicted single-cell data are then extracted by averaging predicted mIF signals within each nucleus, using nuclei masks from an external segmentation model. Finally a cell classifier, trained on validation set cells, predicts cell types. (b) Prediction examples from our best model: Predicted mIF images and cell types (shown as colored cell boundaries) are compared to target mIF images and annotated cell types from the same restained tissue section in the Orion (CD3e, CD8a) and HEMIT (Pan-CK, CD3) datasets. (c) IMMUcan large-area visualization: Nuclei predictions on H&E alongside clustered nuclei from the corresponding consecutive mIF sections.
@@ 5.1.1 Data Configuration
We split the Orion dataset by slide into training (37), validation (2), and test (2) sets. For HEMIT, we used its original train-validation-test splits, while for IMMUcan, all tiles were treated as test set.
All models are trained on H&E-mIF images from ORION extracted as 256x256 pixel tiles at 0.5 mpp. For normalization, target data is scaled to the range $[-0.9,0.9]$ to prevent saturation at extreme values in Tanh activation. Input H&E RGB data is normalized using the mean and standard deviation of the foundation model employed.
To enhance generalization, augmentations include spatial transformations, such as horizontal and vertical flips and coarse dropout (zeroing out a random box in both input and target), along with color augmentations like stain augmentation, random brightness and contrast adjustments, gaussian blur, and gaussian noise. These augmentations simulate variations in stain intensity, color distributions, and common imaging artifacts. To improve robustness, we trained a CycleGAN Zhu et al. (2017) following Runz et al. Runz et al. (2021) to translate Orion H&E images toward the IMMUcan style, capturing more complex domain shift, and used both original and precomputed translated tiles as augmentation during training.
@@ 5.1.2 Training Configuration
We adopt similar Pix2Pix hyperparameters, given its role as a classical image translation model. We use the Adam optimizer with a learning rate of $2\times 10^{-4}$, kept constant in the first half of training before linearly decaying to zero. A linear warmup phase is applied to the generator’s learning rate for the first 400 iterations to improve stability and adaptation, as recommended for ViTs by Dosovitskiy et al. Dosovitskiy et al. (2020). On top of data augmentations, regularization includes weight decay ($1\times 10^{-5}$), gradient clipping (max norm 1), and dropout (0.1). Non-pretrained weights are initialized using a normal distribution with a gain of 0.02, and biases are set to zero. All training are conducted with a batch size of 16. All experiments were performed on a single A100 GPU.
@@ 5.2 Metric Overview
To evaluate our image translation model from H&E to mIF, we used both standard pixel-level and cell-level metrics. At the pixel level, we employed PSNR and SSIM. One of the most important aspect of our model is to capture cell-type information, which is not adequately measured by these metrics. We therefore designed a cell-level fidelity metric.
For this, we extract predicted single-cell data by averaging predicted marker expression within each cell, and classify them using a logistic regression trained on validation set cells. The classifier is then applied to the test set and predictions are compared to pseudo-labels as described in subsection 3.4.4. We use a classifier instead of direct gating on predicted mIF channels to improve robustness to individual channel errors and account for calibration differences between predicted and target fluorescence intensities.
We assess the discriminative power of our model’s predictions by computing Cell AUC (Receiver Operating Characteristic Area Under the Curve), using the predicted mean intensity per nucleus as a score compared to pseudo-labels. To account for class imbalance and better assess model performance, we report F1 score using the cell classifier predictions, as most markers have a lower frequency of positive cells.
We further analyze our model by computing cell count correlations on IMMUcan for CD3, CD8, CD4, FOXP3, and Pan-CK. Since consecutive tissue sections in the IMMUcan dataset are biologically similar but not identical, direct pixel-level alignment is not feasible. However, the minimal spatial separation between sections preserves overall tissue architecture and cellular distribution, allowing for meaningful region-level comparisons. We compute the correlation between predicted and target positive cell counts across registered tiles, using predictions from the logistic regression-based single-cell classification. Following preprocessing in D.1, we compute Pearson correlation coefficients between tile-level cell counts (0.26 mm² area per tile) for each marker. High correlations, even if imperfect, indicate effective cell identification.

### v2

@@ 3.6 Data Preprocessing
Figure 3: Preprocessing pipeline: (a) H&E and mIF images are aligned using Valis (Gatenbee et al., 2023). Tissue regions are then selected via Otsu thresholding on H&E, and a trained CNN filters misaligned tiles caused by the restaining and acquisition process. (b) Autofluorescence is subtracted from mIF images, followed by DAPI-based nuclei segmentation and nuclei dilation to approximate cell boundaries. (c) Pseudo-labels are generated by computing per-cell mean marker expression and applying Gaussian Mixture Model (GMM) clustering to define marker positivity, determining labels (e.g., CD3e+).
We designed a preprocessing pipeline tailored to our H&E-to-mIF prediction task, addressing WSI registration, artifact removal, autofluorescence subtraction, normalization, and cell pseudo-label assignement, as illustrated in Figure 3.
@@ 3.6.1 H&E to mIF registration on consecutive cuts
Accurate registration between H&E and mIF images is crucial for training and evaluating our models. For OrionCRC, HEMIT and PathoCell datasets, data was already registered (Lin et al., 2023a; Bian et al., 2024). For the IMMUcan dataset, we used Valis (Gatenbee et al., 2023) for registration of consecutive H&E and mIF slides.
@@ 3.6.2 Tile Selection
HEMIT and PanNuke are provided as pre-extracted tiles. For OrionCRC, PathoCell, IMMUcan, and Lizard, tiles were extracted from tissue regions identified using Otsu thresholding, with tile sizes defined per dataset (Table 1).
To reduce artifacts from restaining and acquisition in the OrionCRC dataset, we filtered aligned H&E and mIF tiles using several quality control steps. We used thresholding on an empty mIF channel—without associated antibody but capturing shared noise—to identify artifacts affecting all mIF channels. Poor H&E quality tiles were identified by clustering H-optimus-0 embeddings and discarding clusters associated to obvious artifacts. We also manually annotated misaligned tiles and trained a CNN to detect them automatically from H&E and DAPI images. In total, about 40k tiles (10%) were excluded.
For the IMMUcan dataset, where H&E and mIF are from consecutive sections, we selected well-aligned tiles to ensure reliable correlation analysis based on three criteria: Pearson correlation of aligned nuclei density maps from 32×32 patches (threshold=0.25 per tile), tissue overlap ($IoU>0.5$), and tissue percentage ($40\%$). We extracted $1024\times 1024$ pixel tiles at 20x (0.26 mm2), retaining 17k tiles covering 44.4 cm2.
@@ 3.6.3 Autofluorescence Subtraction & Data Normalization
Autofluorescence (AF), captured as an independent channel $I_{AF}$, refers to light naturally emitted by the tissue across channels, introducing noise in other markers. This signal correlates with H&E morphology and can artificially inflate reconstruction metrics if not accounted for. We subtract the AF from each marker channel $I_{IF}^{c}$, using a linear subtraction with parameters $\lambda^{c}$ and $b^{c}$, which were manually adjusted using a Napari (Sofroniew et al., 2025) tool we developed:
$I_{cor}^{c}=\max(0,I_{IF}^{c}-\lambda^{c}\cdot I_{AF}+b^{c})$ (1)
AF subtraction is applied exclusively to the OrionCRC dataset, as it is the only one containing an independent AF channel.
Next, each channel is normalized using the 99.9th percentile $q_{0.999}^{c}$, computed per marker from the distribution of foreground pixel intensities (i.e., values > 0) across the training set, and then log-transformed:
$I_{norm}^{c}=255\cdot\log\left(\frac{\min(I_{cor}^{c},q_{0.999}^{c})}{q_{0.999}^{c}}+1\right)$ (2)
This logarithmic transformation compresses high-intensity values and reduces the impact of extreme outliers, while normalization ensures a consistent dynamic range across markers. This normalization is applied to OrionCRC and PathoCell, while HEMIT mIF is provided already normalized. The autofluorescence Napari tool, selected parameters, code, model, and data access instructions are available on our GitHub, within the Sanofi Public organization, under specific license conditions including a limitation to non-commercial use only.
@@ 3.6.4 Single-Cell Pseudo-Label Extraction
To establish a ground truth for evaluating our model’s ability to identify marker-positive cells from H&E images, we used a standard cell type calling approach on mIF images (Lin et al., 2023b). We first segmented nuclei from the DAPI channel using Cellpose (Stringer et al., 2021) fine-tuned on our data. As a proxy for cell regions, we dilated the nuclear regions by 2 $\mu m$. Single-cell analysis was performed by extracting per-channel mean fluorescence intensities, followed by unsupervised gating using a Gaussian Mixture Model (GMM) to distinguish positive from negative cells (Zhang et al., 2022), with posterior probabilities estimated from the GMM. Although effective, this approach is sensitive to artifacts, approximate boundaries, and signal spillover. To improve robustness, we implemented a hierarchical gating strategy making sure that the biological marker hierarchy was preserved. For instance, given that CD3 positive cells are also CD45 positive, we kept CD3 positivity only for CD45 positive cells. By applying these rules, we obtained a high confidence annotation of our cells.
For the IMMUcan dataset, we used single-cell data provided by the consortium, which was generated using in-house clustering and nucleus segmentation. Since the same nuclei are not necessarily present in consecutive H&E and mIF sections, we applied Hoverfast (Liakopoulos et al., 2024) for nucleus segmentation from H&E images, allowing us to perform the cell-level correlation analysis explained in Section 5.2.
For PathoCell, Lizard, and PanNuke, cell labels were taken directly from the original studies. All three datasets provide nuclei instance masks and cell-type annotations: labels are manually annotated for Lizard and PanNuke, while PathoCell labels are derived from mIF measurements.
Figure 4: MIPHEI architecture: A U-Net-inspired model based on VitMatte for H&E-to-mIF prediction. The main contribution is the integration of a pretrained histology ViT foundation encoder (H-optimus-0) into an image translation model to leverage large-scale pretraining on diverse pathology images. A lightweight Detail Capture module complements the ViT encoder by injecting in the decoder missing spatial information for dense prediction in a U-Net manner. H-Optimus-0 uses a patch size of 14 and patch tokens are resized using bicubic interpolation to match the 1/16 spatial resolution required by the decoder. A Tanh activation is applied, with a custom weighted MSE loss to address marker imbalance. mIF images are shown as composite RGB images (Hoechst in blue, Pan-CK in red).
@@ 4 Methodology
This Section outlines our overall approach for predicting mIF images from H&E images. We used the OrionCRC dataset for training, and HEMIT and IMMUcan datasets for testing.
@@ 4.1 Model Architecture
MIPHEI is a U-Net-inspired architecture designed for image translation from H&E to mIF signals (Figure 4). The framework supports various encoder types, including CNNs like ConvNeXt v2 (Woo et al., 2023), Swin Transformer (Liu et al., 2021), and plain ViTs (Dosovitskiy et al., 2020), enabling integration of recent foundation models in histology, such as CTransPath, Univ2, and H-optimus-0 (Wang et al., 2022; Chen et al., 2024; Saillard et al., 2024). When using CNN-based encoders, MIPHEI follows a standard U-Net formulation, as these encoders naturally provide pyramidal features. For ViT-based encoders, which maintain fixed feature dimensions across layers, MIPHEI adopts a ViTMatte-inspired (Yao et al., 2024) hybrid design: a lightweight convolutional Detail Capture Module extracts multiscale pyramidal features, while ViT features are used as the bottleneck. This module enables the extraction of fine-grained spatial information for dense prediction in a U-Net manner. Unlike original ViTMatte, which concatenates a trimap to the input image, MIPHEI operates directly on the H&E RGB image and omits convolutional necks and window attention mechanisms. Based on the analysis presented in Section 6.1.3, we selected H-optimus-0 as the ViT encoder for MIPHEI.
Building on these encoded features, the decoder reconstructs outputs using bilinear interpolation for upsampling, combined with skip connections, a 3×3 convolutional layer, batch normalization, and ReLU activation. Each mIF marker is predicted by a dedicated output head applied to the decoder output, followed by a Tanh activation.
We also define a MIPHEI variant following UNETR (Hatamizadeh et al., 2022) desing, which employs convolutional transpose operations to upsample ViT token features extracted at different depths. While this approach achieves performance comparable to the ViTMatte-inspired design, the choice of the ViTMatte-based formulation is discussed in Section 6.1.2.
@@ 4.2 Loss Function
We train MIPHEI using a weighted Mean Square Error (MSE) loss to ensure accurate translation from H&E to mIF signals while accounting for varying signal intensities and prevalence across markers. Each marker’s loss is scaled by the inverse of its standard deviation to balance contributions (Mai et al., 2021).
@@
$\mathcal{L}_{\text{wMSE}}=\frac{\lambda}{M}\sum_{j}\frac{1}{\sigma_{j}}L_{\text{MSE},j}$ (3)
@@ 4.3 Foundation Model Training Strategies
While the CNN-based encoders within the U-Net architecture are fully fine-tuned in all experiments, we explore two efficient training strategies for large ViT-based foundation model encoders: (1) decoder-only training with a frozen foundation encoder, leveraging robust pretrained features while reducing trainable parameters; (2) Low-Rank Adaptation (LoRA) (Hu et al., 2021) with rank = 8 and $\alpha=1$, which adapts a pretrained ViT encoder by adding trainable low-rank matrices to the Query and Value projections of the attention layers. Both strategies are common fine-tuning practices for large foundation models when training data is limited to low- or medium-sized datasets.
@@ 5 Experiments
Figure 5: Prediction pipeline and prediction visualization: (a) Inference pipeline: mIF images are first generated using a trained U-Net model. Predicted single-cell data are then extracted by averaging predicted mIF signals within each nucleus, using nuclei masks from an external segmentation model. Finally, a cell classifier, trained on validation set cells, predicts cell types. (b) Prediction examples from our best model: Predicted mIF images and cell types (shown as colored cell boundaries) are compared to target mIF images and annotated cell types from the same restained tissue section in the OrionCRC (CD3e, CD8a) and HEMIT (Pan-CK, CD3) datasets. (c) IMMUcan large-area visualization: Nuclei predictions on H&E alongside clustered nuclei from the corresponding consecutive mIF sections.
@@ 5.1.1 Data Configuration
We split the OrionCRC dataset by slide into training (37), validation (2), and test (2) sets. For HEMIT, we used its original train-validation-test splits. For IMMUcan, Pathocell, Lizard and PanNuke, all tiles were treated as test set.
All models are trained on H&E-mIF images from OrionCRC extracted as 256x256 pixel tiles at 0.5 mpp. For normalization, target data is scaled to the range [0.9,0.9][-0.9,0.9] to prevent saturation at extreme values in Tanh activation. Input H&E RGB data is normalized using the mean and standard deviation of the foundation model employed.
To enhance generalization, augmentations include spatial transformations, such as horizontal and vertical flips and coarse dropout (zeroing out a random box in both input and target), along with color augmentations like stain augmentation, random brightness and contrast adjustments, gaussian blur, and gaussian noise. These augmentations simulate variations in stain intensity, color distributions, and common imaging artifacts. To improve robustness, we trained a CycleGAN (Zhu et al., 2017) following Runz et al. (2021) to translate OrionCRC H&E images toward the IMMUcan style, capturing more complex domain shift, and used both original and precomputed translated tiles as augmentation during training.
@@ 5.1.2 Training Configuration
We adopt similar Pix2Pix hyperparameters (excluding adversarial training), given its role as a classical image translation model. We use the Adam optimizer with a learning rate of $2\times 10^{-4}$, kept constant in the first half of training before linearly decaying to zero. A linear warmup phase is applied to the generator’s learning rate for the first 400 iterations to improve stability and adaptation, as recommended for ViTs by Dosovitskiy et al. (2020). On top of data augmentations, regularization includes weight decay ($1\times 10^{-5}$), gradient clipping (max norm 1), and dropout (0.1). Non-pretrained weights are initialized using a normal distribution with a gain of 0.02, and biases are set to zero. All trainings were conducted with a batch size of 16. All experiments were performed on a single A100 GPU.
@@ 5.2 Metric Overview
To evaluate our models, we report both pixel-level and cell-level metrics.
5.2.1 Pixel-level metrics
At the pixel level, we compute per-channel PSNR, SSIM, and Pearson correlation, and report the average across channels. Pearson correlation is computed over the full evaluation set rather than averaged per tile. As shown in (Balezo et al., 2026), Pearson correlation is less sensitive to background proportion than PSNR and SSIM, making it more suitable for sparse mIF markers.
5.2.2 Cell-level metrics
A key objective of mIF prediction is to preserve biologically meaningful cell-type information for downstream TME analysis. Pixel-level metrics alone do not capture this property, so we also evaluate performance at the cell level. For each nucleus, we extract predicted single-cell expression by averaging the predicted mIF intensity within each cell instance (see Figure 5.a). These predicted cell-level expression vectors are used to train a logistic regression classifier to match cell pseudo-labels (cf Section 3.6.4). The same training and evaluation protocol is applied to all models: for OrionCRC, the cell classifier is trained on validation cells and evaluated on the test set, while for all other external datasets it is trained on a randomly selected 20% subset of cells and evaluated on the remaining 80%. Analogous to linear probing in SSL, this classifier evaluates whether predicted mIF signals preserve discriminative biological structure without relying on complex downstream models. Compared to direct gating on predicted mIF channels, this approach is more robust to individual channel errors and differences in intensity calibration between predicted and real mIF.
Following (Balezo et al., 2026), we report Area Under the Precision-Recall Curve (AUPRC) and F1 score between predicted cell types and cell pseudo-labels on a test set. These metrics are well suited for highly imbalanced cell populations, which are common in mIF-based cell phenotyping, with many more negative than positive cells.
To assess the statistical robustness of our cell-level metrics, we conducted a bootstrap analysis. We generated 1,000 bootstrap samples, each the same size as the original test set, by resampling tiles with replacement. A fixed random seed was used to ensure reproducibility and enable consistent comparison across models. For each sample, we computed the AUPRC and F1 scores and we then reported 95% confidence intervals using the 2.5th and 97.5th percentiles of the resulting distributions. These intervals are presented in Figure 8.
We further analyze our model by computing cell count correlations on IMMUcan for CD3, CD8, CD4, FOXP3, and Pan-CK. Since consecutive tissue sections in the IMMUcan dataset are biologically similar but not identical, direct pixel-level alignment is not feasible. However, the minimal spatial separation between sections preserves overall tissue architecture and cellular distribution, allowing for meaningful region-level comparisons. We compute the correlation between predicted and target positive cell counts across registered tiles, using predictions from the logistic regression-based single-cell classification. Following preprocessing in D.1, we compute Pearson correlation coefficients between tile-level cell counts (0.26 mm2 area per tile) for each marker. High correlations, even if imperfect, indicate effective cell identification.